Could someone please help me with this word problem??? Ed and Dave are going to divide a cash prize of $14,000 in a ratio of 2 to 5. How much will Dave receive?
Answer by checkley75(3666)

2:5
to solve these type of relationships you need to add the 2 factors
2+5=7 so you now have 7 events/items.
now divide the whole by 7
14000/7=2000 for each event/item
seeing as ed gets 2 portions then:
2*2000=4000 answer for Ed.
5*2000=10000 answer for Dave.
proof:
4000:10000 reduces to 2:5